OpenStudy (anonymous):

1.5(2)^x-3=7.6 @lgbasallote

OpenStudy (anonymous):

you need value of x?

OpenStudy (anonymous):

yes!

OpenStudy (lgbasallote):

\[\huge 1.5(2^x) - 3 = 7.6?\]

OpenStudy (anonymous):

ok take log of base two on both sides 1.5log(2^x)-log(3)=log(7.6) or x*1.5log(2)-log(3)=log(7.6) 1.5X=log(7.6)+log(3) use your calculator to solve above them divide by 1.5 and there you are i used calculator and got x=5.64 m i correct ??

OpenStudy (anonymous):

@lgbasallote nop 1.5(2)^x-3=7.6

OpenStudy (lgbasallote):

uhh draw it?

OpenStudy (anonymous):

@sami-21 i think im just going to go with your answer but i got (2)^x= 7.0667

OpenStudy (anonymous):

@lgbasallote u cant draw it

OpenStudy (lgbasallote):

there's a draw button below the reply box

OpenStudy (lgbasallote):

i meant draw the question

OpenStudy (anonymous):